# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

### v1

@@ Abstract
We build upon vec2vec, a procedure designed to align text embedding spaces without parallel data. vec2vec finds a near-perfect alignment, but it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that requires substantially lower computational cost and is highly robust. Moreover, the learned mapping is a linear transformation. Our method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding their results. The method’s stability and interpretable algorithmic steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
@@ 1 Introduction
While the vec2vec approach has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability. GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017). These methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
Our point of departure is still universal geometry, but we pursue a different approach. We draw inspiration from the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations. Unfortunately, stitching requires access to pairs of matching representations. Our key methodological insight is that with universal geometry, structural similarities should be detectable and exploitable, even without adversarial training. Specifically, the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts should provide sufficient signal for alignment, even in the complete absence of parallel supervision. To this end, we utilize the Relative Representation framework of Moschella et al. (2023), which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
@@ 2 Preliminaries
Moschella et al. (2023), inspired by Olah (2015), utilized these observations by proposing relative representations as a mathematically principled framework for comparing embedding spaces. Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable. This work provided both theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
@@ 3 Problem Formulation
We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar semantic structures, they are related by an unknown transformation. Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content. [^1] More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$. We will not take advantage of this flexibility in the paper.
@@ 4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption. Ultimately, the matrix will be the exponentially-smoothed average of a sequence of orthogonal matrices, enforcing orthogonality softly.
@@ 4.2 Approximate Matching
where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices. For solving QAP, we use $2$-OPT (running 30 times and choosing the one with the highest alignment score for further robustness).
@@ 4.4 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima. Pseudo-code is provided in Algorithm 3 & 4.
@@
with $\alpha=0.5$, though the results appear robust to other choices as well.
@@ 5 Experiments
We replicate the in-distribution experiment from Jha et al. (2025). We run the algorithm for all pairs among four sets of text encoders used in Jha et al. (2025): stella, gtr, granite, and e5 (Zhang et al., 2025; Ni et al., 2021; Awasthy et al., 2025; Wang et al., 2024). Experiments use K-Means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT (Croes, 1958).
@@
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. [^3] Cosine similarity in itself is indicative, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity, in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
@@
Note this is a rather economical choice of hyperparameters, which we would not have suggested to use a priori. It turned out to work well, regardless. The full hyperparameter configuration used in the experiments is also documented in the code linked.
@@ Results.
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. Our advantage is more noticeable in the rank metric. Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already at the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $e5$ and gtrgtr pair.
We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise. Our results are not meant to show superiority, but rather matching performance for a fraction of the cost – except in the cases where vec2vec converges poorly, in which we observe mini-vec2vec’s superior behavior. Computationally, one run is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
@@
We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape may contain well-connected regions or multiple basins with similar minima. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
@@ 6 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. Lenc and Vedaldi (2015) and Bansal et al. (2021) demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. Olah (2015) studied the equivalence of neural networks in relative terms. Moschella et al. (2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Huh et al. (2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of usupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018). Most of the works rely on training a GAN. They often require a small seed of matched pairs in order to work. Hoshen and Wolf (2018) avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, specifically their approximate matching process, which relied on PCA, was insufficient for our purposes.
@@ 7 Conclusion
The efficiency and accessibility of our method could democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al. (2025).

### v2

@@ Abstract
We build upon vec2vec, a procedure designed to align text embedding spaces without parallel data. vec2vec finds a near-perfect alignment, but it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that is highly robust and incurs a much lower computational cost. Moreover, the learned mapping is a linear transformation. Our method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding their results. The method’s stability and interpretable algorithmic steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
@@ 1 Introduction
While vec2vec has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability. GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017). These methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
Our point of departure is still universal geometry, but we pursue a different approach. We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations. Unfortunately, stitching requires access to pairs of matching representations. Our key methodological insight is that with universal geometry, structural similarities should be detectable and exploitable, even without adversarial training. Specifically, the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts should provide sufficient signal for alignment, even in the complete absence of parallel supervision. To this end, we utilize the Relative Representation framework of Moschella et al. (2023), which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
@@ 2 Preliminaries
Moschella et al. (2023) operationalized the observations of Olah (2015) and proposed relative representations as a mathematically principled framework for comparing latent spaces. Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable. This work provided both a theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
@@ 3 Problem Formulation
We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar semantic structures, they are related by an unknown transformation. Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content. More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$. We will not take advantage of this flexibility in the paper. [^1] Of course, $f_{1}(x)=f_{2}(x)=\text{const}$ would be a trivial solution, without additional constraints. So, similarity needs to be evaluated in comparison to other vectors in the dataset.
@@
For the gtr-gte pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
@@ 4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption. Ultimately, the matrix will be the exponentially-smoothed average of a sequence of orthogonal matrices, enforcing orthogonality softly.
@@ 4.2 Approximate Matching
where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices. For solving QAP, we use $2$-OPT (Croes, 1958). We run it 30 times and choose the permutation with the highest alignment score. This helps improve the robustness of the alignment over a single run of 2-OPT. Due to the small number of cluster centroids, this takes only a few seconds.
@@ 4.4 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima. Pseudo-code is provided in Algorithms 3 & 4.
@@
In the experiments below, we use $\alpha=0.5$.
@@ 5 Experiments
We replicate the in-distribution experiment from Jha et al. (2025). We run the algorithm for all pairs among the 5 sets of text encoders used in Jha et al. (2025). The encoders are summarized in Table 2. Experiments use k-means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT.
@@
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm. Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry. Our setup does not change the geometry. [^3] Apart from translation and scaling, which, formally speaking, are a form of transition to a third space, but in a very restricted sense.
@@
Note this is a rather economical choice of hyperparameters, which we would not suggest using a priori. It turned out to work well, regardless. The full hyperparameter configuration used in the experiments is also documented in the code linked. For the gtr-gte pair, we increase the number of clusters to $c=30$. We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
@@ Results.
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. The improvement is more noticeable in the rank metric. Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already by the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $e5$-gtrgtr pair or the gtrgtr-gtegte pair.
We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise. The superior results are an additional benefit, but even where the results are not superior, but rather matching performance, it is achieved for a fraction of the cost. Moreover, even where vec2vec collapses, we observe mini-vec2vec still reaches good performance. Computationally, one run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
@@
We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape is well-structured. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
Model Params Backbone Paper Year
gtr 110M T5 Ni et al. 2021
e5 109M BERT Wang et al. 2022
gte 109M BERT Li et al. 2023
stella 109M BERT Zhang et al. 2023
granite 278M RoBERTa Awasthy et al. 2024
Table 2: Embedding models used in our experiments
6 Discussion
Choice of Hyperparameters
The mini-vec2vec algorithm is reasonably stable under different choices of parameters, often converging to similar stationary points. Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and frequently have a monotone behavior. For example, generally speaking, increasing the following hyperparameters should improve the outcome (or at least not harm). Importantly, while some of these trends have been observed empirically in some setups, a full experimental analysis has not been made, and so the explanations below on how to choose hyperparameters are suggestions at best, and should not be trusted blindly.
  • Number of concatenated representations: The more concatenated alignments we use, the more likely it is for noisy assignments to be diluted, emphasizing the signal from true assignments. Note that this helps stabilize the algorithm, meaning it is less likely to meet failure modes in the average and worst cases. However, even with fewer alignments, good alignments can be reached, but the probability of failure modes might increase.
  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows. This makes it a safe bet to increase the number of iterations if one is unsure whether convergence has been reached. Across all cases examined, we found convergence to be quite fast. Almost without exception, 100 iterations were more than enough, by a large margin.
  • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise. The larger the sample size is, the closer this approximation is to the data. We generally consider higher sampling rates to be better.
Other hyperparameters are less obvious. Number of clusters in the approximate matching step is likely good in moderation. While it gives us more landmark points for alignment, it can theoretically lead to instabilities, because the assignment of very granular centroids can be noisy. Increasing number of neighbors is often beneficial, to a point. It helps reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position. Note that this is in contrast to multilingual word embeddings, where essentially every point has a ground-truth translation, and it makes sense to rely on a single neighbor. Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance eventually. Finally, increasing the exponential smoothing parameter $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores. This is likely due to overfitting. The smoothing parameter is implicitly responsible for the maximum distance of the matrix from being orthogonal. While this (alleged) overfitting phenomenon has been observed by us empirically, we did not explore it rigorously, and it might not be a common case. Additionally, increasing $\alpha$ makes convergence slightly slower.
Unlike many deep learning setups – GANs in particular – the range of hyperparameters that lead to strong performance is generously wide. Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter in order to improve effectiveness, if needed.
Failure Modes
We find that despite the many experiments we ran, including ones with different configurations, failure modes were extremely rare. In fact, we didn’t experience any failure modes except with the gte model. Moreover, all failure modes were around the gte-gtr pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in gtr $\rightarrow$ gte – they are just rarer. Crucially, failure modes are not catastrophic. The failure modes are characterized by worse top-1 accuracy and average rank (and slightly lower cosine similarity throughout the refinement steps. Often also noticeable in the initial step, but not always). But the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
@@ 7 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. Lenc and Vedaldi (2015) and Bansal et al. (2021) demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. Olah (2015) studied the equivalence of neural networks in relative terms. Moschella et al. (2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Similarly, Dar et al. (2023) have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors. Huh et al. (2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of unsupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018). Most of the works rely on training a GAN. They often require a small seed of matched pairs in order to work. Hoshen and Wolf (2018) avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, specifically their approximate matching process, which relied on PCA, was insufficient for our purposes.
@@ 8 Conclusion
The efficiency and accessibility of our method could democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al. (2025).
Acknowledgements
We thank John X. Morris for useful feedback. We thank Google for making CPU and GPU notebooks freely available.